OpenStudy (jenniferjuice):

Solve: −5(x − 2) = 15

OpenStudy (anonymous):

divide each side by -5, you'll get x-2=-3

OpenStudy (jenniferjuice):

Solve: −5(x − 2) = 15 1 −3.4 −5 −1

OpenStudy (jenniferjuice):

it doesn't have 3

OpenStudy (anonymous):

i didn't finish it for you

OpenStudy (anonymous):

once you divide by -3, you have x-2 = -3 solve for x

OpenStudy (jenniferjuice):

but wouldn't you have to use the distributive property?

OpenStudy (jenniferjuice):

@callie2240 HELP!

OpenStudy (anonymous):

You can do that, as well but it is easier just to divide by -5. You can do that because you times (x-2) by -5

OpenStudy (jenniferjuice):

yeah i got -5 but im not too sure it is right

OpenStudy (anonymous):

\[-5 * (x-2) = 15\] \[\frac{ -5 * (x-2) }{ -5 }= \frac{ 15 }{ -5 }\] \[(x-2)+2=-3+2\] \[x=-1\]

OpenStudy (amistre64):

you can always dbl chk your results to determine if they are correct or not

OpenStudy (anonymous):

first distribute -5(x-2) that would be -5x+10=15 Now -10 on both sides -5x=5 Now divide by -5 which is x=-1 understand?

OpenStudy (amistre64):

-5( x-2) = 15 -5(-5-2) ?=? 15 -5(-7) ?=? 15 35 ?=? 15 , its not -5

OpenStudy (anonymous):

You can do it either my way, or callie's way. They both work, because the distributive property is a multiplicative property.

OpenStudy (jenniferjuice):

ALRIGHTY! THANKYOU AL SO MUCH <3

OpenStudy (jenniferjuice):

i understand :)